# Improved particle-flow event reconstruction with scalable neural networks for current and future particle detectors

### 29 Jan 2024 v4

@@ Abstract
Experiments at the High-Luminosity LHC and the Future Circular Collider need efficient algorithms to reconstruct granular events expected at such detectors with high fidelity. We study scalable machine learning models for event reconstruction in electron-positron collisions based on a full detector simulation. Particle-flow reconstruction can be formulated as a supervised learning task using tracks and calorimeter clusters. We compare a graph neural network and kernel-based transformer and demonstrate that we can avoid quadratic operations while achieving realistic reconstruction. We show that hyperparameter tuning significantly improves the performance of the models. The best graph neural network model shows improvement in the jet transverse momentum resolution by up to 50% compared to the rule-based algorithm. Accurate reconstruction can significantly improve future measurements at colliders. The resulting model is portable across Nvidia, AMD and Habana hardware. Our datasets and software are published following the findable, accessible, interoperable, and reusable principles.
@@ 1 Introduction
One of the main approaches for event reconstruction at the Large Hadron Collider (LHC) is currently based on the particle-flow (PF) algorithm [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13], which combines measurements from different subdetectors to produce a holistic particle-based description of the entire event. For the planned High Luminosity LHC (HL-LHC) [14] program, due to the installation of new highly granular detector subsystems such as high-granularity calorimeter (HGCAL) for Compact Muon Solenoid (CMS), the reconstruction algorithms have to be revisited or new algorithms have to be developed to fully make use of the data which has a significantly higher complexity. Similarly, for possible future experimental programs such as the Future Circular Collider (FCC) [15, 16], existing algorithms have to be retuned or new ones developed, possibly many times for each new detector scenario under study. Therefore, it is necessary to develop high-fidelity PF reconstruction algorithms that are at the same time computationally efficient, and can be easily extended to new detector concepts without significant manual work. Moreover, if algorithms can be found that can reconstruct granular events with higher fidelity, e.g. in terms of jet response, this may have important implications towards the sensitivity and thus cost-effectiveness of future experiments.
@@
Figure 1: A conceptual overview of the MLPF approach based on tracks, hits and clusters.
@@
We test two types of scalable ML models as benchmarks that can process full events consisting of tens of thousands of measurements, while avoiding memory allocation or computations that scale quadratically with the input size. Using an input of charged particle tracks and clusters of calorimeter energy deposits, we minimize a particle-based loss function and monitor physics metrics that quantify event reconstruction performance during training. We report the results of an extensive hyperparameter optimization (HPO) of the GNN-based model performed using high performance computing (HPC) resources. We then evaluate the model’s physics and computational performance, as well as its portability and quantization compatibility.
Our proposed approach for particle flow reconstruction is summarized in Fig. 1. We find that the GNN-based model can reconstruct physics events with a higher degree of fidelity compared to the rule-based baseline, and its computational cost scales linearly with the size of the input event. Moreover, we demonstrate the portability of the model to several computational hardware processors from Nvidia, AMD, and Intel Habana. The model can also be scaled naturally to lower-level datasets consisting of raw detector hits, if tuned tracking or clustering algorithms are not available. Finally, we identify steps for future development and publish the code and datasets following the findable, accessible, interoperable, and reusable (FAIR) principles [26, 27, 28] for reproducibility and future development.
@@ 2.1 Loss function
The optimization goal for full event particle reconstruction follows the MLPF approach previously used in Refs. [18, 19, 20], and applies a physics-inspired ansatz to simplify the event-based reconstruction loss to a particle-based classification and regression loss.
@@ 2.1.1 Scalable neural network models
However, the evaluation of full self-attention can be computationally demanding and prohibitive for large input multiplicities (${\geq}10^{4}$[32]. Such input multiplicities can be easily reached when considering all tracks and calorimeter clusters in future high luminosity collider events, or if calorimeter clustering is avoided by using tracks and calorimeter hits even in the case of a future lepton collider. It is possible that the local nature of the problem can be exploited directly by first pre-partitioning the event and running the full attention-based model only on subsets. However, this requires careful implementation to avoid artifacts from partitioning and subsequent stitching.
@@
  • The alternative model is a transformer in which the softmax self-attention layer is approximated using positive orthogonal random features [36]. This approach uses a mathematical approximation based on random projections to avoid computing the full attention.
Both of these approaches avoid memory allocations or computations that are quadratic in input multiplicity. The LSH-based approach was first used in [18], initially being paired with a k-nearest neighbors (kNN) graph structure in each bin. In this paper, we instead use an all-to-all fully connected graph in each bin, as we have found it to significantly outperform the kNN based approach computationally as well as in terms of fidelity. Moreover, in this paper, we compare the LSH +GNN approach with the transformer-based alternative systematically. Thus, this paper offers the first detailed comparison of alternative scalable models for particle flow reconstruction.
Figure 2: Top: one layer of the LSH-based GNN. Bottom: one layer of the kernel-based self-attention, figure adapted from [36]. For both layers, the input event of $N$ tracks and clusters (elements) is described by a list of $N$ feature vectors, one $F$-dimensional vector per element. The output of both layers is a list of $N$ transformed feature vectors, one $D$-dimensional vector for each track or cluster. Each input and output vector is represented by a single gray box in the figures. For the LSH-based model, each element is first assigned into a bin based on a learnable function, denoted with the color of the box. The elements are then sorted according to the bin, such that elements in the same bin are consecutive. In each bin, a full all-to-all learnable adjacency matrix is constructed between all the elements in the bin. This matrix is used for message passing in each bin.
@@ 2.1.2 Dataset
Having defined NN models for PF reconstruction that avoid $\mathcal{O}(N^{2})$ memory allocation and computational scaling, we train and test them on a realistic dataset. We generate $\mathrm{e}^{+}\mathrm{e}^{-}$ collision events with Pythia8 (v8.306) [39] and carry out a complete detector simulation with GEANT4 (v11.0.2) using the Key4HEP framework (v2023-01-15) [40]. In particular, we use the Compact Linear Collider (CLIC) detector model [41, 42], along with the Marlin reconstruction code [43], and the Pandora package [44, 45, 46] for a baseline PF implementation. The CLIC detector model is chosen because it is publicly available, well documented and realistic, and similar to detector concepts that are in use at LHC, or under consideration for either HL-LHC or FCC. We note that for HL-LHC, CMS plans the installation of the HGCAL, significantly increasing the event complexity and requiring new approaches for particle reconstruction algorithms, with a view towards balancing high fidelity and computational efficiency, which motivates this study.
@@ 3 Results and discussion
Figure 3: The validation loss, jet and $p_{\mathrm{T}}^{\text{miss}}$ resolution, parameterized by the IQR of the response distribution, for the GNN (red) and transformer-based (blue) models before and after hypertuning, evaluated using 10 trainings, each trained on four Nvidia A100 devices in a distributed data-parallel manner. We show the evolution of these quantities during the full training process, where only the random seed differs in each run. The shaded regions show the standard deviation of the metrics across the runs.
The final model can be trained separately in two different input modes, as visualized in Fig. 1:
  1. (1)
    using charged particle tracks and calorimeter clusters—this tests the physics performance of the model under a typical PF scenario, where a form of clustering is applied beforehand—and
  2. (2)
    using charged particle tracks and raw calorimeter hits—this tests the scalability of the model in cases of high input multiplicity when an explicit clustering step is omitted.
In this paper, we focus on the model (1) based on tracks and clusters.
@@
out_hidden_dims hidden dimension of output decoding MLP
@@
out_hidden_dims hidden dimension of output decoding MLP
@@
An extensive HPO was performed for both the GNN- and transformer-based model using the JURECA supercomputer [52]. JURECA is a pre-exascale modular supercomputer operated by Jülich Supercomputing Centre at Forschungszentrum Jülich. The system consists of a flexible Data Centric (DC) module, based on the Atos BullSequana XH2000. It has, among others, 192 accelerated compute nodes with four NVIDIA A100 GPUs and two AMD EPYC 7742 CPUs each. We employ a similar approach to HPO as in Ref. [53], using Bayesian Optimization in combination with the ASHA [54] algorithm. The optimization was performed in a distributed manner on 96 GPUs spread over 24 compute nodes, consuming roughly 7000 and 5000 GPU hours for the optimization of the transformer and the GNN model respectively. Tables 2 and 2 show the search spaces used for the HPO of the GNN and transformer models respectively, while Tables 4 and 4 show the best HPs found for each model. Both search spaces are relatively large, each containing 10 different HPs with between three and six possible values. Tables 5 and 6 contain descriptions of the HPs in the two search spaces.
The performance improvements achieved from HPO are presented in Fig. 3. We find that the optimized versions of both the GNN and the transformer significantly outperform the unoptimized versions. The optimized version of the GNN significantly outperforms the transformer, although both use a similar number of trainable parameters (${\simeq}5\times 10^{6}$).
The relevant evaluation criterion for the model is the reconstruction of jets and $p_{\mathrm{T}}^{\text{miss}}$ compatible with those at the generator level, a target that is not explicitly trained for, but can be reached by minimizing the particle-level loss. Based on this, we focus on the GNN-based model for the rest of this paper.
@@
Figure 4: Performance of the PF and MLPF algorithms on single particle gun samples. Charged hadrons are shown in the leftmost column, neutral hadrons in the middle column, photons in the right column. Reconstruction efficiency is shown in the top row, fake rate in the middle row, and $p_{\mathrm{T}}$ or $E$ response in the bottom row.
@@
Figure 5: The $p_{\mathrm{T}}$ and $\eta$ distributions of truth particles, baseline PF particles and the particles from the proposed MLPF algorithm for charged hadrons (top row), neutral hadrons (middle) and photons (bottom).
Figure 6: The $p_{\mathrm{T}}$ and $\eta$ distributions of truth particles, baseline PF particles and the particles from the proposed MLPF algorithm for electrons (top row) and muons (bottom row).
Figure 7: The $p_{\mathrm{T}}$ and $\eta$ distributions of jets clustered from truth particles, baseline PF particles and particles from the proposed MLPF algorithm.
@@
Figure 8: Overall jet (top) and total 3-momentum (bottom) response in the $\mathrm{ZH}$, $\mathrm{WW}$ and $\mathrm{t\overline{t}}$ PU10 validation samples that were never used in training.
We use $\mathrm{ZH}$, $\mathrm{WW}$ and $\mathrm{t\overline{t}}$ PU10 events to evaluate out-of-distribution performance. We evaluate the jet response by clustering reconstructed particles into jets as described above, by matching the reconstructed and generator-level jets, and computing the ratio of reconstructed to generator-level jet $p_{\mathrm{T}}$. In all samples, the fraction of reconstructed jets from the GNN-based MLPF is the same or higher than for PF, with generally an improved jet response width, quantified by the IQR and median compared to the rule-based baseline. This improvement over the baseline was not observed before hyperparameter tuning. We also evaluate the total 3-momentum response for either PF or MLPF particles, and find that the MLPF model improves both the median and IQR of the total 3-momentum response distributions for all samples.
@@
Figure 9: The response IQR divided by the median for the jet $p_{\mathrm{T}}$ (left) and total event 3-momentum (right) in the $\mathrm{t\overline{t}}$ PU10 sample. This metric is evaluated in bins of the generator-level jet $p_{\mathrm{T}}$ (left) and generated total event 3-momentum (right).
@@
An important factor in the development of scalable ML-based full event reconstruction models is improving the computational throughput in future deployment scenarios. Possible approaches to improve training throughput could include more efficient model formulations and implementations [56], sparsity and quantization, efficient data loading and distributed training. Therefore, we compare the inference scalability of the rule-based PF implementation on CPU and the proposed scalable GNN implementation on GPU with respect to increasing input multiplicity, with the results summarized in Fig. 10. Inference scaling tests of the GNN model are tested on a small, 8 GB consumer GPU (Nvidia RTX2070S) to emulate a resource-constrained scenario in edge deployment.
@@
Figure 10: Relative timing of the baseline PF (left) and the proposed GNN-based algorithm (right), illustrating the scaling with respect to input particle or element (track, cluster, or hit) multiplicity. The absolute processing time of the baseline algorithm on a single CPU thread is approximately 1,s/event at the reference point of 100 charged pions, which corresponds to approximately $96\pm 3$ tracks and $170\pm 20$ clusters. The runtime of the MLPF algorithm on the RTX2060S at the reference point of $N=256,B=16$ is 2 ms/event. This should not be interpreted as a complete, exhaustive and final computational benchmark, nor should it be used to claim 500-fold improvement in throughput from MLPF, as the absolute timing of any algorithm is heavily dependent on optimizations and hardware. In [20], the runtime of a heavily optimized version of the baseline algorithm was measured at 9 ms/event, whereas the MLPF model at 320 ms/event, both on a single CPU thread.
@@
The rule-based version is run with 25, 50, 100 and 200 generated $\mathrm{\pi^{-}}$ particles per event for 10 events. We generated three different sets of these events varying the random seed each time. The relative runtime of the particle-flow module in Key4HEP is measured, and compared with the 25-particle case. The runtime of the baseline increases nonlinearly with increasing particle multiplicity, and segmentation faults occur for more than approximately 200 particles per event, possibly due to the baseline code and configuration not being tuned for such a high number of particles. Currently, it is only possible to track memory usage of the full baseline reconstruction chain, not individual algorithms. However, we find that the maximum memory requirements increase approximately linearly from about 2 GB for 25 particles to about 8 GB for 200 particles.
@@
The GNN model is run with a varying input size on an 8 GB consumer GPU multiple times to average over random fluctuations. Both the batch size $B$, i.e., the number of batched events ($B\in\{1,2,4,8,16\}$) and the event size $N$, i.e., the number of input elements per event ($N=256n$ for $n\in[1,40]$) are varied independently. The relative runtime with respect to the $N=256,B=16$ case scales approximately linearly with $N$. We note that $B\gg 1$ is required to saturate the GPU, but this is highly model and device specific. The maximum GPU memory of our proposed MLPF algorithm varies between about 300 MB for the smallest tested configurations ($B=1,N=256$ to about 4.5 GB for the largest tested configuration ($B=16,N=10240$), with allocations scaling in a stepwise manner due to the LSH binning.
@@
The training scalability is tested, with results presented in Fig. 11, on three different HPC centers with different accelerator hardware: Nvidia H100 GPUs from Flatiron Institute’s CoreSite cluster [59], AMD MI250 GPUs from the LUMI supercomputer [60], and Intel Habana Gaudi HPUs from the Voyager supercomputer [61]. The LUMI supercomputer features GPU nodes with 64-core AMD Trento CPUs and four AMD MI250X cards, each card consisting of two accelerator chips. Voyager is an NSF-funded supercomputer with 42 first-generation Intel Habana Gaudi (training) nodes, each with eight cards, two first-generation Intel Habana Goya (inference) nodes, a 400 GbE Arista switch, and 3 PB of Ceph file system available at the San Diego Supercomputer Center located at the University of California San Diego. Intel Habana also provided additional access to eight Gaudi2 nodes in an HLS-Gaudi2 Deep Learning Server [62]. For the training tests, there are some differences in the configuration on the HPCs. For the AMD processors, multi-card training was implemented with a mirrored worker configuration, while for the Nvidia and Habana processors, Horovod [63] was used. Events were zero-padded to a regular size of 512 elements per event. A batch size of 250 events per device was used for the Nvidia, AMD, and Habana Gaudi2 processors, while a batch size of 100 per device was used for the Habana Gaudi processors. We observe nearly linear scaling or better for all processors. The scaling for the Habana Gaudi1 (Gaudi2) processors is enhanced by the all-to-all non-blocking intra-node network connection, where each processor has a 100 (300) Gb network connection to every other processor [61, 62].
Figure 11: We demonstrate the scaling of the training performance across multiple devices on a single machine on Nvidia, AMD, and Habana processor cards from the CoreSite, LUMI, and Voyager supercomputers, respectively. For the multi-device scaling test, Horovod [63] was used on CoreSite and Voyager, while the TensorFlow MirroredStrategy [64] was used on LUMI. The batch size was adjusted to fit a single device and the dataset was fully cached in RAM.
We note that when training the track and calorimeter hit based model (2) for approximately ten epochs without any hyperparameter tuning, we observe jet and total 3-momentum response that approaches but does not yet exceed the baseline, motivating further study into efficient training on this dataset for a full end to end reconstruction.

### 08 Mar 2024 v5

@@ Abstract
Efficient and accurate algorithms are necessary to reconstruct particles in the highly granular detectors anticipated at the High-Luminosity Large Hadron Collider and the Future Circular Collider. We study scalable machine learning models for event reconstruction in electron-positron collisions based on a full detector simulation. Particle-flow reconstruction can be formulated as a supervised learning task using tracks and calorimeter clusters. We compare a graph neural network and kernel-based transformer and demonstrate that we can avoid quadratic operations while achieving realistic reconstruction. We show that hyperparameter tuning significantly improves the performance of the models. The best graph neural network model shows improvement in the jet transverse momentum resolution by up to 50% compared to the rule-based algorithm. The resulting model is portable across Nvidia, AMD and Habana hardware. Accurate and fast machine-learning based reconstruction can significantly improve future measurements at colliders.
@@ 1 Introduction
One of the main approaches for event reconstruction at the Large Hadron Collider (LHC) is currently based on the particle-flow (PF) algorithm [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13], which combines measurements from different subdetectors to produce a holistic particle-based description of the entire event. For the planned High Luminosity LHC (HL-LHC) [14] program, due to the installation of new highly granular detector subsystems such as high-granularity calorimeter (HGCAL) for Compact Muon Solenoid (CMS), the reconstruction algorithms have to be revisited or new algorithms have to be developed to fully make use of the data which has a significantly higher complexity. Similarly, for possible future experimental programs such as the Future Circular Collider (FCC) [15, 16], existing algorithms have to be retuned or new ones developed, possibly many times for each new detector scenario under study. Therefore, it is necessary to develop high-fidelity PF reconstruction algorithms that are at the same time computationally efficient, and can be easily extended to new detector concepts without significant manual work. Moreover, if algorithms can be found that can reconstruct events from highly granular detectors with improved fidelity, e.g. in terms of jet response, this may have important implications towards the sensitivity and thus cost-effectiveness of future experiments.
@@
Figure 1: A conceptual overview of the machine-learned particle flow approach based on tracks, hits and clusters on one simulated $\mathrm{t\overline{t}}$ event.
(a) The raw tracker, calorimeter and muon chamber hits, embedded in position space, with the size of the marker proportional to the hit energy. (b) Tracking algorithms reconstruct charged particle tracks from the tracker hits, shown with their extrapolated trajectories. (c) The calorimeter hits are clustered to correspond better to individual particles. (d) The machine-learned particle flow algorithm reconstructs charged and neutral hadrons, photons, electrons and muons based on the tracks and clusters from the previous step, shown with their extrapolated trajectories.
@@
We test two types of scalable ML models as benchmarks that can process full events consisting of tens of thousands of measurements, while avoiding memory allocation or computations that scale quadratically with the input size. Using charged particle tracks and clusters of calorimeter energy deposits, we minimize a particle-based loss function and monitor physics metrics that quantify event reconstruction performance during training. We report the results of an extensive hyperparameter optimization (HPO) of the GNN-based model performed using high performance computing (HPC) resources. We then evaluate the model’s physics and computational performance, as well as its portability and quantization compatibility.
We approach the challenge of high-fidelity full event reconstruction via particle flow using two alternative scalable machine learning models: kernel-based transformers and graph neural networks using locality-sensitive hashing. After a large-scale hyperparameter optimization, we find that the GNN-based model can reconstruct physics events with a higher degree of fidelity compared to the rule-based baseline. At the same time, its computational cost scales linearly with the size of the input event, which is desirable for future deployment scenarios in high-granularity detectors. Moreover, we demonstrate the portability of the model to several computational hardware processors from Nvidia, AMD, and Intel Habana. The model can also be scaled naturally to lower-level datasets consisting of raw detector hits, if tuned tracking or clustering algorithms are not available. Our proposed approach for particle flow reconstruction is summarized in Fig. 1. We identify steps for future development and also publish the code and datasets following the findable, accessible, interoperable, and reusable (FAIR) principles [26, 27, 28] for reproducibility and future development.
@@ 2.1 Loss function
The optimization goal for full event particle reconstruction follows the machine-learned particle flow (MLPF) approach previously used in Refs. [18, 19, 20], and applies a physics-inspired ansatz to simplify the event-based reconstruction loss to a particle-based classification and regression loss.
@@ 2.1.1 Scalable neural network models
However, the evaluation of full self-attention can be computationally demanding and prohibitive for large input multiplicities (${\geq}10^{4}$[32]. Such input multiplicities can be easily reached when considering all tracks and calorimeter clusters in future high luminosity collider events. It is possible that the local nature of the problem can be exploited directly by first pre-partitioning the event and running the full attention-based model only on subsets. However, this requires careful implementation to avoid artifacts from partitioning and subsequent stitching.
@@
  • The alternative model is a kernel-based transformer in which the softmax self-attention layer is approximated using positive orthogonal random features [36]. This approach uses a mathematical approximation based on random projections to avoid computing the full attention.
Both of these approaches avoid memory allocations or computations that are quadratic in input multiplicity. The LSH-based approach was first used in [18], initially being paired with a k-nearest neighbors (kNN) graph structure in each bin. In this paper, we instead use an all-to-all fully connected graph in each bin, as we have found it to significantly outperform the kNN based approach computationally as well as in terms of fidelity. Moreover, in this paper, we compare the LSH +GNN approach with the kernel-based transformer alternative systematically. Thus, this paper offers the first detailed comparison of alternative scalable models for particle flow reconstruction.
Figure 2: One layer of the locality sensitive hashing based graph neural network.
(a) The input event of $N$ tracks and clusters (elements) is described by a list of $N$ feature vectors, one $F$-dimensional vector per element. (b) Each element is first assigned into a bin based on a learnable function, denoted with the color of the box. (c) The elements are then sorted according to the bin, such that elements in the same bin are consecutive. (d) In each bin, a full all-to-all learnable adjacency matrix is constructed between all the elements in the bin, with the learned element-to-element association illustrated by the color of the matrix element. (e) This matrix is used for message passing in each bin, multiplying the corresponding bin feature vectors with the learned adjacency matrix. (f) The output is a list of $N$ transformed feature vectors, one $D$-dimensional vector for each track or cluster. Each input and output vector is represented by a single gray box, the order of the input and output feature vectors is preserved.
@@ 2.1.2 Dataset
Having defined NN models for PF reconstruction that avoid $\mathcal{O}(N^{2})$ memory allocation and computational scaling, we train and test them on a realistic dataset. We generate $\mathrm{e}^{+}\mathrm{e}^{-}$ collision events with Pythia8 (v8.306) [39] and carry out a complete detector simulation with GEANT4 (v11.0.2) using the Key4HEP framework (v2023-01-15) [40]. In particular, we use the Compact Linear Collider (CLIC) detector model [41, 42], along with the Marlin reconstruction code [43], and the Pandora package [44, 45, 46] for a baseline PF implementation. The CLIC detector model is chosen because it is publicly available, well documented and realistic, and similar to detector concepts that are in use at LHC, or under consideration for either HL-LHC or FCC.
@@ 3 Results and discussion
Figure 3: The model validation loss and physics performance throughout training for the graph neural network and kernel-based transformer, before and after hypertuning.
The validation loss (a), jet (b) and $p_{\mathrm{T}}^{\text{miss}}$ resolution (c), parameterized by the interquartile range (IQR) of the response distribution, for the graph neural network (GNN, red) and kernel-based transformer (TF, blue) models before (dashed lines) and after hyperparameter optimization (HPO, solid lines), evaluated using 10 trainings, each trained on four Nvidia A100 devices in a distributed data-parallel manner. Lower values correspond to better physics performance. Only the loss is explicitly minimized, the jet and $p_{\mathrm{T}}^{\text{miss}}$ resolution improvement emerges from the minimization of the loss function. We show the evolution of these quantities during the full training process, where only the random seed differs in each run. The shaded regions show the standard deviation of the metrics across the runs.
An extensive HPO was performed for both the GNN- and kernel-based transformer model using the JURECA supercomputer [52]. JURECA is a pre-exascale modular supercomputer operated by Jülich Supercomputing Centre at Forschungszentrum Jülich. The system consists of a flexible Data Centric (DC) module, based on the Atos BullSequana XH2000. It has, among others, 192 accelerated compute nodes with four NVIDIA A100 GPUs and two AMD EPYC 7742 CPUs each. We employ a similar approach to HPO as in Ref. [53], using Bayesian Optimization in combination with the ASHA [54] algorithm. The optimization was performed in a distributed manner on 96 GPUs spread over 24 compute nodes, consuming roughly 7000 and 5000 GPU hours for the optimization of the kernel-based transformer and the GNN model respectively. In Tables 2 and 2, we show the search spaces used for the HPO of the GNN and kernel-based transformer models respectively, while Tables 4 and 4 show the best HPs found for each model. Both search spaces are relatively large, each containing 10 different HPs with between three and six possible values. Tables 5 and 6 contain descriptions of the HPs in the two search spaces.
@@
out_hidden_dim hidden dimension of output decoding MLP
@@
out_hidden_dim hidden dimension of output decoding MLP
@@
The performance improvements achieved from HPO are presented in Fig. 3. We find that the optimized versions of both the GNN and the kernel-based transformer significantly outperform the unoptimized versions. The relevant evaluation criterion for model selection is the reconstruction of jets and $p_{\mathrm{T}}^{\text{miss}}$ compatible with those at the generator level, a target that is not explicitly trained for, but can be reached by minimizing the particle-level loss. The optimized version of the GNN significantly outperforms the kernel-based transformer, although both use a similar number of trainable parameters (${\simeq}5\times 10^{6}$). As the GNN-based model has significantly improved validation loss and physics performance, we focus on it for the rest of this paper.
@@
Figure 4: Performance of the particle flow (PF) and machine-learned particle flow (MLPF) algorithms on single particle gun samples.
Charged hadron efficiency in (a), neutral hadron efficiency in (b), photon efficiency in (c). The efficiencies are parameterized as a function of track $p_{\mathrm{T}}$ or cluster energy. The charged hadron fake rate is shown in (d), the neutral hadron fake rate in (e), the photon fake rate in (f). For efficiency and fake rate, we show the binomial statistical uncertainties from limited samples.
@@
Figure 5: The generated (truth) and reconstructed kinematic distributions for baseline particle flow (PF) and the proposed machine-learned particle flow (MLPF) algorithm.
The distribution of generated particles for $\mathrm{t\overline{t}}$, $\mathrm{q\overline{q}}$, $\mathrm{WW}$ and $\mathrm{ZH}$, and the ratio between the reconstructed and generated particle distributions for the baseline algorithm (dashed line) and the machine-learned algorithm (solid line) for charged hadrons in (a) and (b), neutral hadrons in (c) and (d), photons in (e) and (f).
Figure 6: The generated (truth) and reconstructed kinematic distributions for baseline particle flow (PF) and the proposed machine-learned particle flow (MLPF) algorithm for electrons and muons.
The distribution of generated particles for $\mathrm{t\overline{t}}$, $\mathrm{q\overline{q}}$, $\mathrm{WW}$ and $\mathrm{ZH}$, and the ratio between the reconstructed and generated particle distributions for the baseline algorithm (dashed line) and the machine-learned algorithm (solid line) for electrons hadrons in (a) and (b) and muons in (c) and (d).
Figure 7: The generated (truth) and reconstructed kinematic distributions for baseline particle flow (PF) and the proposed machine-learned particle flow (MLPF) algorithm for jets.
The $p_{\mathrm{T}}$ (a) and $\eta$ (b) distributions of generated jets for $\mathrm{t\overline{t}}$, $\mathrm{q\overline{q}}$, $\mathrm{WW}$ and $\mathrm{ZH}$, and the ratio between the reconstructed and generated jet distributions for the baseline algorithm (dashed line) and the machine-learned algorithm (solid line).
@@
Figure 8: Jet and total 3-momentum response in the validation samples, comparing the baseline particle flow (PF) and the machine-learned particle flow (MLPF).
The jet response for $\mathrm{ZH}$ samples in (a), for $\mathrm{WW}$ samples in (b) and for $\mathrm{t\overline{t}}$ with hadronic overlay on (c). The total 3-momentum response is shown for $\mathrm{ZH}$ samples in (d), for $\mathrm{WW}$ samples in (e) and for $\mathrm{t\overline{t}}$ with hadronic overlay on (f). We also show the median of the distribution ($M$), for which values closer to unity are better, and the interquartile range (IQR), for which lower values are better.
We use $\mathrm{ZH}$, $\mathrm{WW}$ and $\mathrm{t\overline{t}}$ PU10 events to evaluate out-of-distribution performance. We evaluate the jet response by clustering reconstructed particles into jets, by matching the reconstructed and generator-level jets, and computing the ratio of reconstructed to generator-level jet $p_{\mathrm{T}}$. In all samples, the fraction of reconstructed jets from the GNN-based MLPF is the same or higher than for PF, with generally an improved jet response width, quantified by the interquartile range (IQR) and median compared to the rule-based baseline. This improvement over the baseline was not observed before hyperparameter tuning. We also evaluate the total 3-momentum response for either PF or MLPF particles, and find that the MLPF model improves both the median and IQR of the total 3-momentum response distributions for all samples.
@@
Figure 9: The jet and total 3-momentum resolution, comparing the baseline particle flow (PF) with the machine-learned algorithm (MLPF).
The jet $p_{\mathrm{T}}$ response (a) and total 3-momentum response (b), comparing the baseline particle flow (PF) with the machine-learned algorithm (MLPF), parameterised as the interquartile range (IQR) divided by the median the $\mathrm{t\overline{t}}$ sample with hadronic overlay, evaluated in bins of generated jet $p_{\mathrm{T}}$ (total 3-momentum). Lower values correspond to better resolution.
@@
An important factor in the development of scalable ML-based full event reconstruction models is improving the computational throughput in future deployment scenarios. Possible approaches to improve inference throughput could include more efficient model formulations and implementations [56], sparsity and quantization. Therefore, we compare the inference scalability of the rule-based PF implementation on CPU and the proposed scalable GNN implementation on GPU with respect to increasing input multiplicity, with the results summarized in Fig. 10. Inference scaling tests of the GNN model are tested on a small, 8 GB consumer GPU (Nvidia RTX2070S) to emulate a resource-constrained scenario in edge deployment.
Figure 10: Computational performance of the baseline particle flow (PF) and the proposed machine-learned algorithm in inference and training.
@@
Absolute timing of the baseline PF (a) and the proposed GNN-based algorithm (b), illustrating the scaling with respect to input particle or element (track and cluster) multiplicity. The absolute processing time of the baseline algorithm on a single CPU thread is approximately 1,s/event at the reference point of 100 charged pions, which corresponds to approximately $96\pm 3$ tracks and $170\pm 20$ clusters. The runtime of the machine-learned particle flow (MLPF) algorithm on the Nvidia RTX2060S GPU at the reference point of $N=256$ elements, batch size $B=16$ is 2 ms/event. This should not be interpreted as a complete, exhaustive and final computational benchmark, nor should it be used to claim 500-fold improvement in throughput from MLPF, as the absolute timing of any algorithm is heavily dependent on optimizations and hardware. In [20], the runtime of a heavily optimized version of the baseline algorithm was measured at 9 ms/event, whereas the MLPF model at 320 ms/event, both on a single CPU thread. On panel (c), we demonstrate the scaling of the training performance across multiple devices on a single machine on Nvidia, AMD, and Habana processor cards from the CoreSite, LUMI, and Voyager supercomputers, respectively. For the multi-device scaling test, Horovod [57] was used on CoreSite and Voyager, while the TensorFlow MirroredStrategy [58] was used on LUMI. The batch size was adjusted to fit a single device and the dataset was fully cached in RAM.
@@
The rule-based version is run with 25, 50, 100 and 200 generated $\mathrm{\pi^{-}}$ particles per event for 10 events. We generated three different sets of these events varying the random seed each time, and measured the runtime of the particle-flow module in Key4HEP. The runtime of the baseline increases nonlinearly with increasing particle multiplicity, and segmentation faults occur for more than approximately 200 particles per event, possibly due to the baseline code and configuration not being tuned for such a high number of particles. Currently, it is only possible to track memory usage of the full baseline reconstruction chain, not individual algorithms. However, we find that the maximum memory requirements increase approximately linearly from about 2 GB for 25 particles to about 8 GB for 200 particles.
@@
The GNN model is run with a varying input size on an 8 GB consumer GPU multiple times to average over random fluctuations. Both the batch size $B$, i.e., the number of batched events ($B\in\{1,2,4,8,16\}$) and the event size $N$, i.e., the number of input elements per event ($N=256n$ for $n\in[1,40]$) are varied independently. The inference runtime scales approximately linearly with $N$. We note that $B\gg 1$ is required to saturate the GPU, but this is highly model and device specific. The maximum GPU memory of our proposed MLPF algorithm varies between about 300 MB for the smallest tested configurations ($B=1,N=256$ to about 4.5 GB for the largest tested configuration ($B=16,N=10240$), with allocations scaling in a stepwise manner due to the LSH binning.
@@
The training scalability is also tested, with results presented in Fig. 10, on three different HPC centers with different accelerator hardware: Nvidia H100 GPUs from Flatiron Institute’s CoreSite cluster [61], AMD MI250 GPUs from the LUMI supercomputer [62], and Intel Habana Gaudi HPUs from the Voyager supercomputer [63]. The LUMI supercomputer features GPU nodes with 64-core AMD Trento CPUs and four AMD MI250X cards, each card consisting of two accelerator chips. Voyager is an NSF-funded supercomputer with 42 first-generation Intel Habana Gaudi (training) nodes, each with eight cards, two first-generation Intel Habana Goya (inference) nodes, a 400 GbE Arista switch, and 3 PB of Ceph file system available at the San Diego Supercomputer Center located at the University of California San Diego. Intel Habana also provided additional access to eight Gaudi2 nodes in an HLS-Gaudi2 Deep Learning Server [64]. For the training tests, there are some differences in the configuration on the HPCs. For the AMD processors, multi-card training was implemented with a mirrored worker configuration, while for the Nvidia and Habana processors, Horovod [57] was used. Events were zero-padded to a regular size of 512 elements per event. A batch size of 250 events per device was used for the Nvidia, AMD, and Habana Gaudi2 processors, while a batch size of 100 per device was used for the Habana Gaudi processors. We observe nearly linear scaling or better for all processors. The scaling for the Habana Gaudi1 (Gaudi2) processors is enhanced by the all-to-all non-blocking intra-node network connection, where each processor has a 100 (300) Gb network connection to every other processor [63, 64].